OpenStudy (lisa2345):

I need help b^-3= 1/216 can someone teach me to answer this?

OpenStudy (campbell_st):

well you need to know about negative powers... \[a^{-b} = \frac{1}{a^b}\] so in your question its \[\frac{1}{b^3} = \frac{1}{216}\] if you cross multiply, or take the reciprocals of each fraction you'll be able to solve for b. hope it helps

OpenStudy (anonymous):

1/(6^3) = 1/216

OpenStudy (lisa2345):

thankyou guys so much :}